OpenStudy (anonymous):

Write the series in summation notation, using the symbol ∑ and appropriate indices, Do NOT find the indicated sum. 9+13+17+21+25 (I need to find the top and bottom number of ∑)

OpenStudy (anonymous):

first term = 9 difference = +4 Tn = a +(n-1)d = 9 +4(n-1) = 4n +5 so its the sum of (4n+5) from n=1 to 5

OpenStudy (anonymous):

\[\sum_{n=1}^{5} ( 4n+5) \]

OpenStudy (anonymous):

wow u guys are incredible... thank you...!!!